Presentation is loading. Please wait.

Presentation is loading. Please wait.

Status of the CTP O.Villalobos Baillie University of Birmingham April 23rd 2009.

Similar presentations

Presentation on theme: "Status of the CTP O.Villalobos Baillie University of Birmingham April 23rd 2009."— Presentation transcript:

1 Status of the CTP O.Villalobos Baillie University of Birmingham April 23rd 2009

2 Plan of Talk Introduction Hardware development Firmware development Software development Plans for next few months.

3 Starting Point Last September, the CTP was ready for the first collisions, and made considerable progress over the short period with circulating beams, even managing alignment of some trigger inputs In November, the CTP was switched off, under the condition that, exceptionally, it could be operated at specific request from a detector during the shutdown period. There have been a few requests, e.g. from TPC, Muon Trigger What else has happened?

4 Development Programme Hardware Improve optical connection to beam data devices BST1, BST2 in order to extract both clock and data from both Install optical fibre and drivers to transmit scope signals from trigger boards to surface Firmware LO board. Downscaling by dead time. LTU Maximum rate cut (requested by TPC) Upgrade to Switch firmware Software Improvements to offline interface, monitoring software and additions to make use of upgrades described above.

5 Hardware Installation of required fibres for BST1, BST2 readout In progress. BST2 signal required in two places so needs to be split. Suitable splitter being identified. Transmission of scope signals In progress. Delays with manufacturer in supplying suitably fast optical driver. Is now supplying such driver, but only with wrong optical connector. In order to make progress, adaptor is needed for use with our fibres. This would allow prototype to be made. Schedule now quite tight for having final version installed in time for collisions.

6 Firmware CTP upgrade New mode of downscaling. Programmable class trigger BUSY time. LTU upgrade. Done. (Rate control plus some minor improvements, mainly affecting those few detectors still using TTCvi.) Switch upgrade. Done. Now needs new software to exploit new features. See below.

7 Trigger Tests at Point2 status last year

8 Software 1 CTP Counters data readout Currently under review at trigger meetings organized by Federico Antinori. (see presentation by Anton Jusko April 22nd 2009) 6 counters per class (counts before and after every trigger level applied) saved in OCDB. Needed by offline for cross section calculations. All CTP counters saved in DCS for monitoring purposes – kept for “several months”. 8 values kept per LTU in DCS, again for monitoring purposes. Interaction records are sent to DAQ by DDL for each orbit. DAQ currently stores IRs only for orbits associated with L2a triggers (+/- 2 orbits). What to do with rest of data under discussion.

9 Software 2 Orbit-Turn checks CTP has an orbit counter (24 bits) that increments by one every time the requisite number of BCs is counted. “Turn” counter available on RF2TTC board, which notes orbit changes as seen by machine (using 32 bit counters) Anton has implemented program to look for jumps in difference between counters. Results indicate difference is stable with TTCmi– does not increase, but can fluctuate by +/- 1 turn Almost certainly not worrying as comparison relies on a network transfer that does not have a definite time. (average 60 microseconds) Increase seen with BST message –under investigation

10 Software 3 Switch upgrade allows counter readout for all inputs whether in CTP or not. Will be read out and stored to DCS Comparison with CTP counts for relevant counters will be made.

11 Plans for Next few months For more details see talk by Anton Jusko at run coordination meeting next Tuesday. CTP ON again. May-June Repeat of all commissioning tests with individual detectors as done last year (in conjunction with DAQ). Before August. CTP/ECS/DAQ integration: CTP readout and interaction record; advanced configurations (multiple clusters and classes) Stabilize transfer of.rcfg files to OCDB; recheck synchronization of LTUs; rehearse trigger alignment procedure. Set up calibration triggers Implement required p/f protections for different triggers. August Official cosmic data run September Start of injection tests. Start using alignment tools.

12 Summary Some important upgrades have been made during shutdown period, which should make very useful improvements to CTP functionality. With possible exception of oscilloscope transmission (dependent on receipt of components) all should be completed well in time. Programme for startup of trigger well advanced. Will be presented on Tuesday.

13 L1 Spurious Monitoring L1 gated so L1 spurious suppressed New counters introduced on L1 and fanout boards counting glitches before gatin g. Hints at correlation between glitches and VME activity.

Download ppt "Status of the CTP O.Villalobos Baillie University of Birmingham April 23rd 2009."

Similar presentations

Ads by Google